# On De Giorgi’s Conjecture of Nonlocal approximations for free-discontinuity problems: The symmetric gradient case

### v1

@@ 3.1. Preliminaries
For every $u\in\textnormal{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ it holds $\mathscr{I}_{u,1}(B)=\zeta_{1}(B)$ for every Borel set $B\subset\Omega$.
@@
Let $u\colon\Omega\to\mathbb{R}^{n}$ be measurable and assume that there exists $p\in(1,\infty]$ such that $\mathscr{I}_{u,p}$ is finite. Then,
$\mathscr{I}_{u,1}\big{(}\{x\in\Omega\colon\text{$0\notin J_{\hat{u}^{\xi}_{x}}$ for $\mathcal{H}^{n-1}$-a.e.~{}$\xi\in\mathbb{S}^{n-1}$}\}\big{)}=0\,.$ (3.15)
@@
Notice that $\hat{u}^{\xi}_{x}(t)=\hat{u}^{\xi}_{\pi_{\xi}(x)}(x\cdot\xi+t)$ for every $t\in\Omega^{\xi}_{x}$, whenever $\pi_{\xi}\colon\mathbb{R}^{n}\to\Pi^{\xi}$ denotes the orthogonal projection. By letting $t_{x}:=x\cdot\xi$, then the set in (3.15) coincides with
$\{x\in\Omega\colon t_{x}\notin J_{\hat{u}^{\xi}_{\pi_{\xi}(x)}}\text{ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$}\}.$ (3.16)
The proof follows exactly as in [2, Proposition 4.7]. ∎
@@
The proposition above is a direct consequence of Riesz Theorem. To keep the presentation self contained, we include a proof below.
Proof.
[…]
By renormalization, with no loss of generality, we may assume that for every $i=1,\dotsc,k$ and every $j=1,2,\dotsc$ it holds true $|\alpha^{j}_{i}|\leq 1$ and there exists at least one $i(j)=1,\dotsc,k$ such that $|\alpha^{j}_{i(j)}|=1$. Up to pass to a not relabelled subsequence in $j$, we may suppose that the index $i(j)$ does not depend on $j$, and that $\alpha^{j}_{i}\to\alpha_{i}$ for every $i=1,\dotsc,n$, for some real coefficients $\{\alpha_{1},\dotsc,\alpha_{k}\}$ which are not all identically zero. Therefore, since $K_{j}\nearrow A$ in measure we deduce that $\sum_{i=1}^{k}\alpha^{j}_{i}i_{K_{j}}(v_{i})$ (extended to zero out of $K_{j}$) converges to $\sum_{i=1}^{k}\alpha_{i}v_{i}$ pointwise a.e. in $A$ as $j\to\infty$. From (3.17) we immediately deduce that $\sum_{i=1}^{k}\alpha_{i}v_{i}=0$ in $L^{0}(A)$ which gives a contradiction. The claim is thus proved.
Recall the definition of $\Theta^{*n-1}$ from [24].
@@
Let $u\in{\rm GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$. Then, for $\hat{\mu}^{1}_{u}$-a.e. $x\in\Omega$, the condition $\Theta^{*n-1}(\hat{\mu}^{1}_{u},x)=0$ implies that the set $\{\xi\in\mathbb{S}^{n-1}:0\in J_{\hat{u}^{\xi}_{x}}\}$ is $\mathcal{H}^{n-1}$-negligible.
@@
$\Sigma:=\{\xi\in\mathbb{S}^{n-1}:0\in J_{\hat{u}^{\xi}_{x}}\}$
@@
Now define for every $0<r<\text{dist}(x,\partial\Omega)$ the rescaled function $u_{r}\colon B_{1}(0)\to\mathbb{R}^{n}$ as $u_{r}(z):=u(x+rz)$ and define the truncated cone
$\text{co}(\Sigma\cup\{0\}):=\{z=t\xi:\xi\in\Sigma,\ t\in(-1,1)\setminus\{0\}\}.$
From formula (3.3), we deduce that $u_{r}\in\text{GBV}^{\mathcal{E}}(B_{1}(0);\mathbb{R}^{n})$ and that $\mu^{\xi}_{u_{r}}(B_{1}(0))=r^{1-n}\mu^{\xi}_{u}(B_{r}(x))$ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$. Let now $\{e_{1},\dotsc,e_{n}\}$ denote the orthonormal basis of $\mathbb{R}^{n}$, and consider the sets $\Lambda^{\pm}:=\{\xi\in\mathbb{S}^{n-1}:\pm(e_{1}\cdot\xi)>0\}$. We notice that $\mathbb{S}^{n-1}\setminus(\Lambda^{+}\cup\Lambda^{-})=\mathbb{S}^{n-1}\cap\Pi^{e_{1}}$, $\Lambda^{+}\cap\Lambda^{-}=\emptyset$, and $\Lambda^{-}=-\Lambda^{+}$. Let us further denote the two maps $\phi^{\pm}\colon\mathbb{R}^{n}\setminus\Pi^{e_{1}}\to\Lambda^{\pm}$ defined as
$\phi^{\pm}(z):=\begin{cases}\frac{z}{|z|},&\text{ if }\frac{z}{|z|}\in\Lambda^{\pm}\\ -\frac{z}{|z|},&\text{ if }\frac{z}{|z|}\in\Lambda^{\mp}.\end{cases}$
Notice that, by construction, the functions $t\mapsto\phi^{+}(t\xi)$ and $t\mapsto\phi^{-}(t\xi)$ are constants where they are defined, namely, for $\xi\in\mathbb{S}^{n-1}\setminus\Pi^{e_{1}}$ and $t\in\mathbb{R}\setminus\{0\}$, and hence they can be continuously extended to the whole of $\mathbb{R}$.
Since $\mathcal{H}^{n-1}(\Sigma)>0$ and since $\mathbb{S}^{n-1}\cap\Pi^{e_{1}}$ is a negligible set, then $\mathcal{H}^{n-1}(\Sigma\cap\Lambda^{+})>0$ or $\mathcal{H}^{n-1}(\Sigma\cap\Lambda^{-})>0$. Hence, we may assume with no loss of generality that $\Sigma$ is all contained for instance in $\Lambda^{+}$. The case in which $\Sigma$ is all contained in $\Lambda^{-}$ can be treated exactly in the same way. From our assumption, we have $\mathcal{L}^{n}(\text{co}(\Sigma\cup\{0\}))>0$ and for $\xi\in\Sigma$ we obtain
$(\hat{u}_{r})^{\xi}_{0}\to f_{\xi}\ \text{ in }\ L^{1}((-1,1))\text{ as $r\to 0^{+}$, \ where }\ f_{\xi}(t):=a_{\xi}\,\text{sign}(t)+b_{\xi}$ (3.21)
for some $a_{\xi}\in\mathbb{R}\setminus\{0\}$ and $b_{\xi}\in\mathbb{R}$. Condition (3.21) tells us in particular that $u_{r}\cdot\phi^{+}\to f$ in measure on $\text{co}(\Sigma\cup\{0\})$ as $r\to 0^{+}$, where $f\colon B_{1}(0)\to\mathbb{R}$ is defined as
$f(t\xi):=\begin{cases}f_{\xi},&\text{ if $t\xi\in\text{co}(\Sigma\cup\{0\})$ for $(\xi,t)\in\Lambda^{+}\times\mathbb{R}$},\\ 0,&\text{ otherwise}.\end{cases}$ (3.22)
Step 3.
We claim that there exist a subsequence of radii $r_{k}\searrow 0$ and linearly independent vectors $\{z_{1},\dotsc,z_{n}\}\subset B_{1}(0)\setminus\{0\}$, such that $z_{i}$ are points of approximate continuity for $u_{r_{k}}$, while (3.18) holds true for $x=z_{i}$ and $u=u_{r_{k}}$, and $O_{z_{i}}(u_{r_{k}})\to 0^{+}$ for every $k=1,2,\dotsc$ as $k\to\infty$. Notice that, once the claim is proved, by the Fundamental Theorem of Calculus together with a measure theoretic argument we find that for $\mathcal{L}^{n}$-a.e. $z\in B_{1}(0)$ and for every $1=1,\dotsc,n$
$|(u_{r_{k}}(z_{i})-u_{r_{k}}(z))\cdot\phi^{+}(z-z_{i})|$ $\leq|D(\hat{u}_{r})_{z_{i}}^{\phi^{+}(z-z_{i})}|(B_{1}(0)_{z_{i}}^{\phi^{+}(z-z_{i})}),\ k=1,2,\dotsc$ (3.23)
We notice that the right-hand side of (3.23) is $\mathcal{L}^{n}$-measurable as a function of $z$. Indeed, we could have chosen the vectors $\{z_{1},\dotsc,z_{n}\}$ so that $\xi\mapsto|D(\hat{u}_{r})_{z_{i}}^{\xi}|(B_{1}(0)_{z_{i}}^{\xi})$ are measurable maps between the $\sigma$-algebra of $\mathcal{H}^{n-1}$-measurable subsets of $\mathbb{S}^{n-1}$ and the Borel $\sigma$-algebra of $\mathbb{R}$ (this follows similarly as in the proof of Lemma 2.1). Then, $z\mapsto|D(\hat{u}_{r})_{z_{i}}^{\phi^{+}(z-z_{i})}|(B_{1}(0)_{z_{i}}^{\phi^{+}(z-z_{i})})$ is simply the composition of $\xi\mapsto|D(\hat{u}_{r})_{z_{i}}^{\xi}|(B_{1}(0)_{z_{i}}^{\xi})$ with $z\mapsto\phi^{+}(z-z_{i})$, which is Borel measurable and whose preimages of $\mathcal{H}^{n-1}$-negligible sets are $\mathcal{L}^{n}$-negligible.
To prove the claim we first notice that the assumption $\Theta^{n-1}(\hat{\mu}^{1}_{u},x)=0$ implies that $\hat{\mu}^{1}_{u_{r}}(B_{1}(0))\to 0$ as $r\to 0^{+}$. By letting $\lambda_{r}:=\hat{\mu}^{1}_{u_{r}}(B_{1}(0))$ we see from (3.20) applied to $u=u_{r}$ and $\Omega=B_{1}(0)$ that
$\mathcal{L}^{n}(\{z\in B_{1}(0):O_{z}(u_{r})>c_{n}(\Omega)\sqrt{\lambda_{r}}\})\leq\sqrt{\lambda_{r}},\ \text{ for every $0<r<\text{dist}(x,\partial\Omega)$}.$
Therefore, by passing to a subsequence $r_{k}\searrow 0$ such that $\sum_{k}\sqrt{\lambda_{r_{k}}}\leq\mathcal{L}^{n}(B_{1}(0))/2$, we have that the set $A:=\bigcap_{k}\{z\in B_{1}(0):O_{z}(u_{r_{k}})\leq c_{n}(\Omega)\sqrt{\lambda_{r_{k}}}\}$ satisfies $\mathcal{L}^{n}(A)\geq\mathcal{L}^{n}(B_{1}(0))/2$. In particular, since $\mathcal{L}^{n}(A)>0$, we can find $\{z_{1},\dotsc,z_{n}\}\subset A$ in generic position such that $z_{i}$ is an approximate continuity point for $u_{r_{k}}$ and (3.18) holds true with $u=u_{r_{k}}$ and $x=z_{i}$ for every $k=1,2,\dotsc$ and $i=1,\dotsc,n$. The fact that $O_{z_{i}}(u_{r_{k}})\to 0^{+}$ for every $i=1,\dotsc,n$ as $k\to\infty$ is a direct consequence of the fact that $\{z_{1},\dotsc,z_{n}\}\subset A$. The claim is thus proved.
Step 4.
Condition $O_{z_{i}}(u_{r_{k}})\to 0^{+}$ as $k\to\infty$ allows us to say that, up to considering a not relabelled subsequence,
$g_{i,k}(\xi)\to 0\ \ \text{ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ \ \ as \ \ $k\to\infty$,}$ (3.24)
where we have set
$g_{i,k}(\xi):=|D(\hat{u}_{r_{k}})^{\xi}_{z_{i}}|(B_{1}(0)^{\xi}_{z_{i}}\setminus J^{1}_{(\hat{u}_{r_{k}})^{\xi}_{z_{i}}})+\mathcal{H}^{0}(B_{1}(0)^{\xi}_{z_{i}}\cap J^{1}_{(\hat{u}_{r_{k}})^{\xi}_{z_{i}}}).$
In particular, for every sufficiently large $k$ (depending on $\xi$) we have that for every $i=1,\dotsc,n$
$g_{i,k}(\xi)=|D(\hat{u}_{r_{k}})^{\xi}_{z_{i}}|(B_{1}(0)^{\xi}_{z_{i}}).$ (3.25)
Now since $\mathcal{L}^{n}(\text{co}(\Sigma\cup\{0\}))>0$ and since $\{z_{1},\dotsc,z_{n}\}$ are linearly independent vectors, we obtain
[…]

### v2

@@ 3.1. Preliminaries
For every $u\in\textnormal{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ and for every Borel set $B\subset\Omega$, it holds
$\mathscr{I}_{u,1}(B)=\int_{\mathcal{S}^{n-1}}\eta_{\xi}(B)\textnormal{d}\mathcal{H}^{n-1}(\xi)=\int_{\mathcal{S}^{n-1}}\Big{(}\int_{\Pi^{\xi}}\sum_{t\in B^{\xi}_{y}}\big{(}|[\hat{u}^{\xi}_{y}(t)]|\wedge 1\big{)}\,\mathrm{d}\mathcal{H}^{n-1}(y)\Big{)}\textnormal{d}\mathcal{H}^{n-1}(\xi).$
Given $E\subset\mathbb{R}^{n}$, we define the codimension one slice of $E$ as
$E^{\xi}_{s}:=\{y\in\Pi^{\xi}+s\xi:y\in E\}.$
Proposition 3.11.
Let $\xi\in\mathbb{S}^{n-1}$, $s\in\mathbb{R}$ and let $u\in\textnormal{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$. Then for $\mathcal{H}^{n-1}$-a.e. $y\in\Omega_{s}^{\xi}$ there exists
$\operatorname*{ap-\lim}_{\begin{subarray}{c}z\to y,\\ z\in(H^{\pm}_{\xi}+s\xi)\cap\Omega\end{subarray}}u(z)=:u^{\xi\pm}_{s}(y),$ (3.15)
where $H^{+}_{\xi}:=\{z\in\mathbb{R}^{n}:z\cdot\xi>0\}$ and $H^{-}_{\xi}$ is defined similarly.
Proof.
Since $u\in\textnormal{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$ we can find a basis $\{\xi_{1},\ldots,\xi_{n}\}$ of $\mathbb{R}^{n}$ such that $\hat{\mu}_{u}^{\xi_{i}}(\Omega)<\infty$ for $i=1,\ldots,n$. By applying [26, Theorem 5.1], we infer that
$\operatorname*{ap-\lim}_{\begin{subarray}{c}z\to y,\\ z\in(H^{\pm}_{\xi_{i}}+s\xi_{i})\cap\Omega\end{subarray}}u(z)\cdot\xi_{i}=:v_{i}^{\pm}(y)$
exists for every $i=1,\ldots,n$ and for $\mathcal{H}^{n-1}$-a.e. $y\in\Omega_{s}^{\xi}$. Since $\xi_{1},\ldots,\xi_{n}$ form a basis we can readily conclude that the limit in (3.15) exists and, in particular, $u^{\xi\pm}_{s}(y)\cdot\xi_{i}=v_{i}^{\pm}(y)$. ∎
Thanks to Proposition 3.11, we are now in a position to identify a precise representative for co-dimension one slices through $(n-1)$-planes.
Definition 3.12.
Given $\xi\in\mathbb{S}^{n-1}$ and $u\colon\Omega\to\mathbb{R}^{n}$, we define for every $s\in\mathbb{R}$ the function $\bar{u}^{\xi}_{s}\colon\Pi^{\xi}+s\xi\to\Pi^{\xi}$ as
$\bar{u}^{\xi}_{s}(y):=\pi_{\xi}(u_{s}^{\xi+}(y)),\quad y\in\Omega^{\xi}_{s},$
where $u_{s}^{\xi+}$ is defined in (3.15).
Proposition 3.13.
Let $u\in\textnormal{GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$. Then for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ and for $\mathcal{L}^{1}$-a.e. $s\in\mathbb{R}$ the function $\bar{u}^{\xi}_{s}\in\textnormal{GBV}^{\mathcal{E}}(\Omega_{s}^{\xi};\Pi^{\xi}).$
Proof.
[…]
By using (2.2) and (3.16) it holds for $\mathcal{H}^{n-1}$-a. $\xi\in\mathbb{S}^{n-1}$
$\infty$ $>\int_{\Pi^{\xi}\cap\mathbb{S}^{n-1}}\hat{\mu}_{u}^{\eta}(\Omega)\textnormal{d}\mathcal{H}^{n-2}(\eta)$
$=\int_{\Pi^{\xi}\cap\mathbb{S}^{n-1}}\Big{(}\int_{\Pi^{\eta}}(\hat{\mu}_{u})_{y}^{\eta}(\Omega_{y}^{\eta})\textnormal{d}\mathcal{H}^{n-1}(y)\Big{)}\textnormal{d}\mathcal{H}^{n-2}(\eta).$
[…]
$\int_{\Pi^{\xi}\cap\mathbb{S}^{n-1}}\Big{(}\int_{\Pi^{\eta}}(\hat{\mu}_{u})_{y}^{\eta}\textnormal{d}\mathcal{H}^{n-1}(y)\Big{)}\textnormal{d}\mathcal{H}^{n-2}(\eta)$
$=\int_{\Pi^{\xi}\cap\mathbb{S}^{n-1}}\Big{(}\int_{\mathbb{R}}\Big{(}\int_{\Pi^{\eta}\cap(\Pi^{\xi}+s\xi)}(\hat{\mu}_{u})_{y}^{\eta}(\Omega_{y}^{\eta})\textnormal{d}\mathcal{H}^{n-2}(y)\Big{)}\textnormal{d}\mathcal{L}^{1}(s)\Big{)}\textnormal{d}\mathcal{H}^{n-2}(\eta).$
Finally we observe that, for every $\xi,\,\eta\in\mathbb{S}^{n-1}$, $u\cdot\eta=\pi_{\xi}(u^{+})\cdot\eta$ $\mathcal{L}^{n}$-a.e., thus we can substitute in the above integral $(\hat{\mu}_{u})_{y}^{\eta}=(\hat{\mu}_{\bar{u}_{s}^{\xi}})_{y}^{\eta}$, which concludes the proof. ∎
@@
Assume that Theorem 3.6 holds true up to dimension $n-1$. Let $u\colon\Omega\to\mathbb{R}^{n}$ be measurable. Then, it holds
$\mathscr{I}_{u,1}\big{(}\{x\in\Omega:\text{for every }\xi\in\mathbb{S}^{n-1},x\notin J_{\bar{u}_{s}^{\xi}}\text{ with }s\textnormal{ s.t. }x\in\Pi^{\xi}+s\xi\}\big{)}=0\,.$ (3.17)
@@
Let $E$ be the set in (3.17), i.e.,
$E:=\{x\in\Omega:\text{for every }\xi\in\mathbb{S}^{n-1},x\notin J_{\bar{u}_{s}^{\xi}}\text{ with }s\textnormal{ s.t. }x\in\Pi^{\xi}+s\xi\}.$ (3.18)
We show that $E$ is $\hat{\mu}_{u}^{1}$-measurable. We proceed as in [2, Lemma 4.4]. Let $\tau:=\arctan$, we define $s^{\pm}:\Omega\times\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\to\mathbb{R}^{n}$ and $i^{\pm}:\Omega\times\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\to\mathbb{R}^{n}$ as
$s_{j}^{\pm}(x,\xi,\nu):=\limsup_{r\searrow 0}\int_{(\Pi^{\xi}+(x\cdot\xi)\xi)\cap(H_{\nu}^{\pm}+x)\cap B_{r}(x)}\tau((\bar{u}^{\xi}_{x\cdot\xi})_{j}(z))\,\textnormal{d}\mathcal{H}^{n-1}(z),$
$i_{j}^{\pm}(x,\xi,\nu):=\liminf_{r\searrow 0}\int_{(\Pi^{\xi}+(x\cdot\xi)\xi)\cap(H_{\nu}^{\pm}+x)\cap B_{r}(x)}\tau((\bar{u}^{\xi}_{x\cdot\xi})_{j}(z))\,\textnormal{d}\mathcal{H}^{n-1}(z),$
for $j=1,\ldots,n$. We notice that $s^{\pm}$ and $i^{\pm}$ are Borel measurable in $\Omega\times\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}$. Indeed, the integrand functions are Borel measurable in the variables $(x,\xi,\nu,z)$. Hence, Fubini’s theorem implies that the integral functions are Borel measurable in $\Omega\times\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}$. Finally, both liminf and limsup can be computed by restricting $r\in\mathbb{Q}$ because of the continuity of the integrals with respect to $r$. Then, the set
$B=\left\{(x,\xi,\nu)\in\Omega\times\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}:s^{+}(x,\xi,\nu)=i^{+}(x,\xi,\nu),\ s^{-}(x,\xi,\nu)=i^{-}(x,\xi,\nu),\right.$
$\left.s^{+}(x,\xi,\nu)\neq s^{-}(x,\xi,\nu),\quad s^{\pm}_{j}(x,\xi,\nu)\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)\text{ for $j=1,\ldots,n$}\right\},$
is Borel measurable. Furthermore, $\Omega\setminus E=\pi_{1}(B)$, where $\pi_{1}\colon\Omega\times\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\to\Omega$ is the projection over the first component. Then, by the measurable projection Theorem (see for instance [29, Section 2.2.13]) the set $\Omega\setminus E$ is $\hat{\mu}_{u}^{1}$-measurable, and the same holds for $E$.
From the very definition of $E$, we note that it holds $(E\cap(\Pi^{\xi}+s\xi))\cap J_{\bar{u}_{s}^{\xi}}=\emptyset$ for every $\xi\in\mathbb{S}^{n-1}$ and $s\in\mathbb{R}$. Moreover, we know from Proposition 3.13 that for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ and for $\mathcal{L}^{1}$-a.e $s\in\mathbb{R}$ we also have $\bar{u}_{s}^{\xi}\in\textnormal{GBV}^{\mathcal{E}}(\Omega_{s}^{\xi};\Pi^{\xi})$.
For simplicity let us denote $v:=\bar{u}_{s}^{\xi}$. Then, by the assumption that Theorem 3.6 holds true, we have, for $\mathcal{L}^{1}$-a.e. $s\in\mathbb{R}$,
$J_{\hat{v}_{y}^{\eta}}=(J_{v})_{y}^{\eta}$ (3.19)
for $\mathcal{H}^{n-2}$-a.e. $\eta\in\mathbb{S}^{n-1}\cap\Pi^{\xi}$ and $\mathcal{H}^{n-2}$-a.e. $y\in\Pi^{\eta}\cap(\Pi^{\xi}+s\xi)$. Since the map $(z,s)\mapsto u^{\xi}_{s}(z)$ coincides with the map $u(z)$ for $\mathcal{H}^{n-1}$-a.e. $z\in\Omega_{s}^{\xi}$ and for $\mathcal{L}^{1}$-a.e. $s\in\mathbb{R}$, and since $\eta\in\Pi^{\xi}$, we have as well
$u(y+t\eta)\cdot\eta=\pi_{\xi}(u^{\xi}_{s}(y+t\eta))\cdot\eta=v(y+t\eta)\cdot\eta,$
for $\mathcal{L}^{1}$-a.e. $s\in\mathbb{R}$, for $\mathcal{H}^{n-2}$-a.e. $y\in\Pi^{\eta}\cap(\Pi^{\xi}+s\xi)$, and for $\mathcal{L}^{1}$-a.e. $t\in\Omega^{\eta}_{y}$. In particular, by virtue of Fubini’s Theorem (up to measurability issues….), and using also (3.19), we infer that for $\mathcal{H}^{n-2}$-a.e. $\eta\in\mathbb{S}^{n-1}\cap\Pi^{\xi}$ it holds
$J_{\hat{u}_{y}^{\eta}}\cap E_{y}^{\eta}=J_{\hat{v}_{y}^{\eta}}\cap E_{y}^{\eta}=(J_{v})^{\eta}_{y}\cap E^{\eta}_{y}=(J_{v}\cap E)^{\eta}_{y}=\emptyset\quad\text{for }\mathcal{H}^{n-1}\text{-a.e. }y\in\Pi^{\eta},$
where in the last equality we have used the relation $(E\cap(\Pi^{\xi}+s\xi))\cap J_{v}=\emptyset$ valid for every $s\in\mathbb{R}$. Summarizing we have obtained that for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$
$J_{\hat{u}_{y}^{\eta}}\cap E_{y}^{\eta}=\emptyset,\ \ \text{ for $\mathcal{H}^{n-2}$-a.e. $\eta\in\mathbb{S}^{n-1}\cap\Pi^{\xi}$ and $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\eta}$}.$ (3.20)
Eventually, by (3.16), we infer from (3.20) that
$J_{\hat{u}_{y}^{\xi}}\cap E_{y}^{\xi}=\emptyset,\ \ \text{ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$ and $\mathcal{H}^{n-1}$-a.e. $y\in\Pi^{\xi}$}.$
From the very definition of the measures $\eta_{\xi}$ (cf. (3.13)) we obtain that $\eta_{\xi}(E)=0$ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$. As a direct consequence of the definition of $\mathscr{I}_{u,1}$ we conclude $\mathscr{I}_{u,1}(E)=0$. ∎
@@
The proposition above is a direct consequence of Riesz Theorem (see for instance [37, Theorem 1.21]). To keep the presentation self contained, we include a proof in the Appendix A.
Recall the definition of $\Theta^{*n-1}$ (see [29] for further details)
$\Theta^{*n-1}(\mu,x):=\limsup_{r\to 0}\frac{\mu(B_{r}(x))}{r^{n-1}},$
for every measure $\mu\in\mathcal{M}^{+}(\Omega)$ and $x\in\Omega$.
@@
Let $u\in{\rm GBV}^{\mathcal{E}}(\Omega;\mathbb{R}^{n})$. Then, for $\hat{\mu}^{1}_{u}$-a.e. $x\in\Omega$, the condition $\Theta^{*n-1}(\hat{\mu}^{1}_{u},x)=0$ implies that $x\notin J_{\bar{u}_{s}^{\xi}}$ for every $\xi\in\mathbb{S}^{n-1}$, with $s\in\mathbb{R}$ such that $x\in\Pi^{\xi}+s\xi$.
@@
$\Sigma:=\{\xi\in\mathbb{S}^{n-1}:x\in J_{\bar{u}^{\xi}_{s}}\}$
is not empty. Fix $\xi\in\Sigma$. By the same arguments used to prove the measurability of the set $E$ defined in (3.18), it can be shown that $\Sigma$ is $\mathcal{H}^{n-1}$-measurable.
Let $x\in\Omega$, then $x\in\Pi^{\xi}+s\xi$ with $s=x\cdot\xi$. Now define for every $0<r<\text{dist}(x,\partial\Omega)$ the rescaled function $u_{r}\colon B_{1}(0)\to\mathbb{R}^{n}$ as $u_{r}(z):=u^{+}(x+rz)$, and $\bar{u}_{r}\colon B_{1}(0)\cap\Pi^{\xi}\to\mathbb{R}^{n}$ as $\bar{u}_{r}(z):=\pi_{\xi}(u^{+}(x+rz))$.
From formula (3.3), we deduce that $u_{r}\in\text{GBV}^{\mathcal{E}}(B_{1}(0);\mathbb{R}^{n})$ and that $\hat{\mu}^{\xi}_{u_{r}}(B_{1}(0))=r^{1-n}\hat{\mu}^{\xi}_{u}(B_{r}(x))$ for $\mathcal{H}^{n-1}$-a.e. $\xi\in\mathbb{S}^{n-1}$. Let us further denote the map $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined as
$\phi(z):=\begin{cases}\frac{z}{|z|},&\text{ if }z\neq 0\\ 0,&\text{ if }z=0.\end{cases}$
From our assumption, we have, for $\xi\in\Sigma$, that there exists $\nu\in\Pi^{\xi}\cap\mathbb{S}^{n-1}$ such that
$\bar{u}_{r}\to f_{\nu,\xi}\ \text{ in measure }\text{ as $r\to 0^{+}$, \ where}\ f_{\nu,\xi}(z):=a_{\nu,\xi}\,\text{sign}(z\cdot\nu)+b_{\nu,\xi}$ (3.24)
for some $a_{\nu,\xi}\in\mathbb{R}^{n}\setminus\{0\}$ and $b_{\nu,\xi}\in\mathbb{R}^{n}$.
Step 3.
We claim that there exist a subsequence of radii $r_{k}\searrow 0$ and vectors $\{z_{1},\dotsc,z_{n2^{n}}\}\subset B_{1}(0)\setminus\{0\}$, such that $\{z_{(j-1)n+1},\ldots,z_{jn}\}$ are linearly independent vectors contained in the $j$-th orthant of $\mathbb{R}^{n}$, $z_{i}$ are points of approximate continuity for $u_{r_{k}}$, while (3.21) holds true for $x=z_{i}$ and $u=u_{r_{k}}$, and $O_{z_{i}}(u_{r_{k}})\to 0^{+}$ for every $k=1,2,\dotsc$ as $k\to\infty$. Notice that, once the claim is proved, by the Fundamental Theorem of Calculus together with the concavity of the truncation function, there exists $j=1,\ldots,2^{n}$ such that for $\mathcal{H}^{n-1}$-a.e. $z\in B_{1}(0)\cap\Pi^{\xi}$ and for every $i=(j-1)n+1,\dotsc,jn$, we have $\xi\cdot z_{i}>0$ and
$|(u_{r_{k}}(z_{i})-u_{r_{k}}(z))\cdot\phi(z-z_{i})|\wedge 1$ $\leq|\hat{\mu}_{{u}_{r}}^{\phi(z-z_{i})}|(B_{1}(0)),\ k=1,2,\dotsc$ (3.25)
We notice that the right-hand side of (3.25) is $\mathcal{H}^{n-1}$-measurable as a function of $z$. Indeed, we could have chosen the vectors $\{z_{1},\dotsc,z_{n2^{n}}\}$ so that $\eta\mapsto|\hat{\mu}_{{u}_{r}}^{\eta}|(B_{1}(0))$ are measurable maps between the $\sigma$-algebra of $\mathcal{H}^{n-1}$-measurable subsets of $\mathbb{S}^{n-1}$ and the Borel $\sigma$-algebra of $\mathbb{R}$ (this follows similarly as in the proof of Lemma 2.1). Then, $z\mapsto|\hat{\mu}_{u_{r}}^{\phi(z-z_{i})}|(B_{1}(0))$ is simply the composition of $\eta\mapsto|\hat{\mu}_{u_{r}}^{\eta}|(B_{1}(0))$ with $z\mapsto\phi(z-z_{i})$, which is Borel measurable and whose preimages of $\mathcal{H}^{n-1}$-negligible sets are $\mathcal{H}^{n-1}$-negligible.
To prove the claim we first notice that the assumption $\Theta^{n-1}(\hat{\mu}^{1}_{u},x)=0$ implies that $\hat{\mu}^{1}_{u_{r}}(B_{1}(0))\to 0$ as $r\to 0^{+}$. By letting $\lambda_{r}:=\hat{\mu}^{1}_{u_{r}}(B_{1}(0))$ we see from (3.23) applied to $u=u_{r}$ and $\Omega=B_{1}(0)\cap\Sigma_{j}$, where $\Sigma_{j}$ is the $j$-th orthant, that
$\mathcal{L}^{n}(\{z\in B_{1}(0)\cap\Sigma_{j}:O_{z}(u_{r})>c_{n}(\Omega)\sqrt{\lambda_{r}}\})\leq\sqrt{\lambda_{r}},\ \text{ for every $0<r<\text{dist}(x,\partial\Omega)$}.$
Therefore, by passing to a subsequence $r_{k}\searrow 0$ such that $\sum_{k}\sqrt{\lambda_{r_{k}}}\leq\mathcal{L}^{n}(B_{1}(0))/2^{n+1}$, we have that the set $A:=\bigcap_{k}\{z\in B_{1}(0)\cap\Sigma_{j}:O_{z}(u_{r_{k}})\leq c_{n}(\Omega)\sqrt{\lambda_{r_{k}}}\}$ satisfies $\mathcal{L}^{n}(A)>0$. In particular, we can find $\{z_{(j-1)n+1},\dotsc,z_{jn}\}\subset A\cap\Sigma_{j}$ in generic position such that $z_{i}$ is an approximate continuity point for $u_{r_{k}}$ and (3.21) holds true with $u=u_{r_{k}}$ and $x=z_{i}$ for every $k=1,2,\dotsc$ and $i=(j-1)n+1,\dotsc,jn$. The fact that $O_{z_{i}}(u_{r_{k}})\to 0^{+}$ for every $i=(j-1)n+1,\dotsc,jn$ as $k\to\infty$ is a direct consequence of the fact that $\{z_{(j-1)n+1},\dotsc,z_{jn}\}\subset A$. The claim is thus proved.
Step 4.
Fix $j=1,\ldots,2^{n}$ such that $\xi\cdot z_{i}>0$ for $i=(j-1)n,\ldots,jn$. For simplicity assume $j=1$. We observe that for every $i=1,\ldots,n$
$\lim_{r\to 0}\hat{\mu}_{u_{r}}^{\phi^{+}(z-z_{i})}(B_{1}(0))=0\qquad\text{for $\mathcal{H}^{n-1}$-a.e.~{}$z\in\Pi^{\xi}\cap B_{1}(0)$}$ (3.26)
since $\Theta^{*n-1}(\hat{\mu}^{1}_{u},x)=0$.
Since $\{z_{1},\dotsc,z_{n}\}$ are linearly independent vectors, we obtain that